# AT2023lli: A Tidal Disruption Event with Prominent Optical Early Bump and Delayed Episodic X-ray Emission

### v1

@@ Abstract
High-cadence, multiwavelength observations have continuously revealed the diversity of tidal disruption events (TDEs), thus greatly advancing our knowledge and understanding of TDEs. In this work, we conducted an intensive optical-UV and X-ray follow-up campaign of TDE AT2023lli, and found a remarkable month-long bump in its UV/optical light curve nearly two months prior to maximum brightness. The bump represents the longest separation time from the main peak among known TDEs to date. The main UV/optical outburst declines as , making it one of the fastest decaying optically selected TDEs. Furthermore, we detected sporadic X-ray emission 30 days after the UV/optical peak, accompanied by a reduction in the period of inactivity. It is proposed that the UV/optical bump could be caused by the self-intersection of the stream debris, whereas the primary peak is generated by the reprocessed emission of the accretion process. In addition, our results suggest that episodic X-ray radiation during the initial phase of decline may be due to the patched obscurer surrounding the accretion disk, a phenomenon associated with the inhomogeneous reprocessing process. The double TDE scenario, in which two stars are disrupted in sequence, is also a possible explanation for producing the observed early bump and main peak. We anticipate that the multicolor light curves of TDEs, especially in the very early stages, and the underlying physics can be better understood in the near future with the assistance of dedicated surveys such as the deep high-cadence survey of the 2.5-meter Wide Field Survey Telescope (WFST).
@@ 1 Introduction
A tidal disruption event (TDE) occurs when a star gets too close to the supermassive black hole (SMBH) at the center of a galaxy and is torn apart by its gravitational force. In recent years, many TDEs have been detected by sky survey projects using various telescopes and instruments (Gezari, 2021; Hammerstein et al., 2023; Jiang et al., 2021; van Velzen et al., 2019, 2021; Wang et al., 2022; Yao et al., 2023; Zhu et al., 2023). These TDEs commonly show a blue continuum that can be characterized as a black body of nearly constant temperature () over time, broad emission lines (such as H and He ii) and a “fast-rise and slow-decay” pattern (Arcavi et al., 2014; van Velzen et al., 2020; Gezari, 2021).
@@
Most TDEs that have been discovered so far exhibit a smooth single-peak light curve in both UV and optical bands. However, some of them display a bump in the rising phase of their light curves (Holoien et al., 2019; Charalampopoulos et al., 2023; Faris et al., 2023; Wang et al., 2023b; Huang et al., 2023b, Lin et al. in prep.). This early signature may be produced by a number of physical processes, such as the cooling of unbounded debris, vertical compression during the first pericenter passage, and shock breakout of self-crossing debris (Kasen & Ramirez-Ruiz, 2010; Yalinewich et al., 2019; Wang et al., 2023b). But no definitive conclusion has been reached. The collision between stream debris and the disk was proposed to explain the early bump in recent UV/optical outbursts of the candidate for repeated partial TDE ASASSN-14ko (Huang et al., 2023b). Furthermore, a fraction of the sources appeared to rebrighten in the optical light curve as the duration of observation increased (Jiang et al., 2019; Malyali et al., 2021; Chen et al., 2022; Wang et al., 2023c; Yao et al., 2023). Rebrightening in TDE has also been found in X-rays, and the likely source is repeated partial TDE (Liu et al., 2023, 2024; Miniutti et al., 2023; Wevers et al., 2023). Repeated partial TDE has been applied to explain the periodic optical outburst of ASASSN-14ko (Payne et al., 2021) and the optical rebrightening in AT2020vdq (Somalwar et al., 2023).
@@
The source of the X-ray emission is unquestionably the accretion processes, but there is an ongoing debate regarding the origin of the UV/optical emissions. Two rival models are being discussed: one suggests that UV/optical emission is a result of the reprocessing of X-ray photons in an extended envelope/outflow (Loeb & Ulmer, 1997; Strubbe & Quataert, 2009; Metzger & Stone, 2016; Roth et al., 2016; Dai et al., 2018), while the other proposes that the emissions are caused by shocks generated by the stream-stream collision of the stellar debris (Piran et al., 2015; Jiang et al., 2016; Huang et al., 2023c; Jankovič et al., 2023). An intriguing phenomenon in optical TDEs is that X-ray emission lags UV/optical emission (Gezari et al., 2017; Pasham et al., 2017; Liu et al., 2022; Chen et al., 2022; van Velzen et al., 2021; Shu et al., 2020; Wang et al., 2022, 2023c; Huang et al., 2023a). This can be explained by a structural change in the reprocessing layer as the accretion rate declines from super-Eddington to sub-Eddington (e.g. Thomsen et al. 2022) or by a delayed onset of accretion after circularization.
@@
AT2023lli (RA=22:57:39.470, DEC=+40:32:40.02) is a TDE discovered by the Gravitational-wave Optical Transient Observer (GOTO) on 2023 Jun 23, at a redshift of 0.036111https://www.wis-tns.org/object/2023lli. We conducted multiwavelength follow-up observations of this source and found that it showed a strong UV/optical bump and a delayed and episodic X-ray emission. In Section 2, we describe the observations and data analysis methods, and the multiwavelength results are represented in Section 3. In Section 4, we discuss the possible physical mechanisms for the early bump and the X-ray emission. In Section 5, we summarize our main results and conclusions. In this work, we assume cosmological parameters of , , and .
@@ 2.1 Swift X-ray Photometry
We proposed several target of opportunity (ToO) observations to monitor the multiwavelength evolution using the X-ray telescope (XRT; Burrows et al. 2005) and the ultraviolet/optical telescope (UVOT; Roming et al. 2005) onboard the Swift Observatory. We also retrieved the public data (PI: Hinkle) from the High Energy Astrophysics Science Archive Research Center website. Swift data were processed with HEASoft 6.32.1. We ran the tasks xrtpipeline and xrtproducts to generate light curves and spectra. The source and background count rates were extracted from circular regions centered on the object with a radius of and a nearby circular region with a radius of , respectively. Source detection was performed using the ximage task. For data in which the source was not detected, we derived a 3 upper limit on the X-ray flux using the WebPIMMS tool222https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl and assuming a power-law X-ray spectrum with a photon index of 1.75 (Ricci et al., 2017). To improve the signal-to-noise ratio, we stacked the event files for MJD 60198-60201 (ObsIDs: 00016100029 and 00016100030) and MJD 60226-60244 (ObsIDs: 00016100038, 00016100039, and 00016100040), respectively. In these two stacked files, the source was detected above the 3 level using the ximage task.
@@ 2.2 XMM-Newton Observation
We proposed a ToO observation with X-ray Multi-Mirror Mission (XMM-Newton), which was performed on MJD 60295.5 (ObsID: 0932391301) with a total observation time of 20 ks. The XMM-Newton data were reduced with the Science Analysis System (SAS version 21.0.0). Tasks cifbuild and odfingest were executed for data preparation, and then light curves and spectra were extracted through the task xmmextractor.
@@ 2.4 Ground-based Optical Photometry
AT2023lli was first detected by the Asteroid Terrestrial Impact Last Alert System (ATLAS; Tonry et al. 2018; Smith et al. 2020). The ATLAS - and -band light curves were obtained using the ATLAS Forced Photometry Service, which produces point spread function (PSF) photometry on the difference images. ATLAS has 3–4 single exposures within each epoch (typically within one day), so we binned the light curve every epoch to improve the signal-to-noise ratio (SNR).
Upon noticing that AT2023lli displayed a rebrightening after the initial peak (the “early bump”), we promptly conducted optical monitoring in the bands using 1.0 m telescopes from the Las Cumbres Observatory Global Telescope network (LCOGT; Brown et al. 2013) starting from July 28, 2023. The observation continued until October 18, when it almost faded out in the LCOGT exposures. The observations were carried out every 2–3 days with a denser sampling around the peak, using exposure times of 260s, 160s, and 260s in the , , and bands, respectively.
Subsequently, we also initiated a follow-up campaign with the 2.5 m Wide Field Survey Telescope (WFST) in the bands with an exposure time of 30 s from 2023 September 14. There is a gap between November 21 and December 29, due to engineering maintenance and testing. The exposure time was changed to 60 s after the observations were resumed. Briefly, the WFST is a newly established photometric facility equipped with a 2.5-meter diameter primary mirror, an active optics system, and a mosaic CCD camera with 0.73 gigapixels on the primary focal plane, allowing high-quality image capture over a 6.5 square-degree field of view (Wang et al., 2023a). It was installed near the summit of Saishiteng Mountain in northwestern China in the summer of 2023 and entered its commissioning stage in September following the official release of the first-light image of the Andromeda Galaxy333https://wfst.ustc.edu.cn/news-and-meetings/news/20230917/20230917/. Follow-up observations with the WFST are particularly important when AT2023lli becomes fainter, making its detection challenging with 1m LCOGT telescopes.
@@ 2.5 Spectral Observations and Data Reduction
Shortly after its discovery, the SCAT group acquired a spectrum of AT2023lli utilizing the Supernova Integral Field Spectrograph (SNIFS), which was installed on the UH88 telescope (Hinkle 2023). This spectrum was retrieved from the TNS website555https://www.wis-tns.org/object/2023lli. We then acquired four more spectra using the Double Spectrograph (DBSP) instrument mounted on the Hale 200-inch telescope at Palomar Observatory (P200, Oke & Gunn 1982). For these observations, the D55 dichroic was adopted to split the incoming light longer than or shorter than 5500 Å into separate red and blue channels. A grism of 600 lines per mm blazed at 3780 Å was used for the blue arm, and a grism of 316 lines per mm blazed at 7150 Å for the red arm. The data were then reduced using the Python package Pypeit (Prochaska et al. 2020a, b), which can fully automatically implement the standard reduction procedure for long-slit spectroscopic observations. Furthermore, we obtained another spectrum using the Yunnan-Faint Object Spectrograph and Camera (YFOSC) on board the LiJiang 2.4m telescope (Fan et al. 2015; Wang et al. 2019) on January 2, 2024. All spectra are shown in Figure 3. Broad (10,000 km s) and probable Helium lines are detected when the source is bright. As the TDE emission significantly diminished, the spectrum (+147 and + 1365 days) is dominated by host galaxy emission, showing deep Balmer absorption (see Figure 3), consistent with the types of galaxies that TDEs are known to preferentially inhabit (French et al., 2016).
@@ 3.1 Multiwavelength Light Curves
The multiwavelength light curves of AT2023lli show a complex structure with an early bump and a main peak in UV/optical bands. The source was first detected by GOTO on MJD 60118, with a magnitude of 18.84 mag in the L band (Hinkle, 2023). We followed up the initial observation by ATLAS and found that the source had already been visible on MJD 60112.57 with a magnitude of mag in the c band. The source was then brightened to mag in the same band by MJD 60116.56, before fading. We missed the early UV emission of the source. The first UVOT observation was made on MJD 60125.5, when the source was already declining, with host-subtracted magnitudes of , and mag for the UVW1, UVM2 and UVW2 bands, respectively. The source then rose again after MJD 60140 and reached the optical peak on MJD 60164, with the magnitudes of , and mag for the g, r, i bands, respectively. At that time, we obtained the magnitudes of , , , and mag for the UVW1, UVM2, UVW2 and U bands, respectively.
@@
We stacked a total of 27 Swift/XRT event files with an effective exposure time of 59.1 ks, covering the period from MJD 60125 to 60188. From these stacked data, we obtained a 3 upper limit of 0.0011 corresponding to a luminosity of by fixing the Galactic hydrogen column density value of (HI4PI Collaboration et al., 2016) and assuming a photon index of 1.75 (Ricci et al., 2017). The source was initially detected in the X-ray band until MJD 60198, where a 2.5 detection was achieved with a count rate of . Additionally, the source was also detected on MJD 60200. To improve the signal-to-noise ratio, we combined the event files from two ObsIDs (00016100029 and 00016100030), which covered the time period from MJD 60198 to 60200. From these stacked data, we derived a mean count rate of at the 3 level, with an effective exposure time of 4013 s. However, in a subsequent observation with Swift/XRT lasting 4152 s, only a 3 upper limit of 0.004 was obtained. By merging three event files from MJD 60226 to 60237 (ObsIDs: 00016100038–00016100040), we derived a mean count rate of . In the following Swift/XRT observation with an exposure time of 3678 s, the source was not detected, and a 3 upper limit of 0.003 is imposed. Thus, the X-ray emission from the source appeared to exhibit a sporadic behavior with a general upward trend. At MJD 60295.5, XMM-Newton recorded an unabsorbed luminosity of in the X-rays. Following this, the X-ray emission was detectable in every subsequent observation with XRT. When the X-ray showed sporadic behavior, we checked the data and found that the count rate was above the statistical fluctuation and confirmed that such behavior actually existed (P-value of 0.0004). For XRT, the conversion factor from count rate to flux was determined to be , which was derived from the spectral fit of the stacked spectrum extracted from all event files with X-ray detection.
@@
The UV/optical SEDs were fitted with a black body model and SuperBol (Nicholl, 2018) was used to calculate the temporal evolution of the black body luminosity (), temperature () and black body radius (). The results are displayed in Figure 4. The highest was recorded on MJD 60164.6 with a value of and at that time we derived and . After that, luminosity decreases, while no significant temporal evolution can be seen in the black body temperature. We derived the energy released in UV/optical bands of and during the bump and the main outburst, respectively. Using the emcee python package (Foreman-Mackey et al., 2013), we employed a “Gaussian rise and power-law decay” model to fit in both the bump and the main outburst, as described in van Velzen et al. (2019, 2021); Hammerstein et al. (2023); Yao et al. (2023). We obtained a power-law index of -4.10, which is much steeper than the theoretical prediction of -5/3 (Rees, 1988). Moreover, some researchers found that TDEs exhibit a wide range of power-law indices, from -4.0 to -0.5 (van Velzen et al., 2019, 2021; Gezari, 2021; Hammerstein et al., 2023). Our result resembles that of AT2019mha, the most rapidly fading source (with a power-law index of 4.0) among the optically-selected TDEs studied by van Velzen et al. (2021). This makes AT2023lli one of the fastest declining TDEs ever observed. A steep slope in the light curve has been predicted by some theoretical work (Guillochon & Ramirez-Ruiz, 2013; Ryu et al., 2020) that involves partial TDEs (pTDEs). As showed by Ryu et al. (2020), a more rapid decrease in light curves is associated with a lower amount of mass being removed from the star. Consequently, the situation of AT2023lli could be connected to a pTDE. A recent systematic analysis conducted on optically-selected TDEs has found that the ratio between UV/optical and X-ray luminosities () tends to approach values similar to those of a disk () at later times, although it exhibits a wide range of values at earlier times (Guolo et al., 2023). This trend is also observed in the case of AT2023lli. As the X-ray emission becomes visible and increases in brightness, the ratio gradually increases, with the most recent value slightly above 1 (refer to the bottom panel of Figure 4).

### v2

@@ Abstract
High-cadence, multiwavelength observations have continuously revealed the diversity of tidal disruption events (TDEs), thus greatly advancing our knowledge and understanding of TDEs. In this work, we conducted an intensive optical-UV and X-ray follow-up campaign of TDE AT 2023lli and found a remarkable month-long bump in its UV/optical light curve nearly two months prior to maximum brightness. The bump represents the longest separation time from the main peak among known TDEs to date. The main UV/optical outburst declines as , making it one of the fastest- decaying optically selected TDEs. Furthermore, we detected sporadic X-ray emission 30 days after the UV/optical peak, accompanied by a reduction in the period of inactivity. It is proposed that the UV/optical bump could be caused by the self-intersection of the stream debris, whereas the primary peak is generated by the reprocessed emission of the accretion process. In addition, our results suggest that episodic X-ray radiation during the initial phase of decline may be due to the patched obscurer surrounding the accretion disk, a phenomenon associated with the inhomogeneous reprocessing process. The double TDE scenario, in which two stars are disrupted in sequence, is also a possible explanation for producing the observed early bump and main peak. We anticipate that the multicolor light curves of TDEs, especially in the very early stages, and the underlying physics can be better understood in the near future with the assistance of dedicated surveys such as the deep high-cadence survey of the 2.5 m Wide Field Survey Telescope.
@@ 1 Introduction
A tidal disruption event (TDE) occurs when a star gets too close to the supermassive black hole (SMBH) at the center of a galaxy and is torn apart by its gravitational force. In recent years, many TDEs have been detected by sky survey projects using various telescopes and instruments (van Velzen et al., 2019, 2021; Gezari, 2021; Jiang et al., 2021; Wang et al., 2022; Hammerstein et al., 2023; Yao et al., 2023; Zhu et al., 2023). These TDEs commonly show a blue continuum that can be characterized as a blackbody of nearly constant temperature () over time, broad emission lines (such as H and He ii) and a “fast rise and slow decay” pattern (Arcavi et al., 2014; van Velzen et al., 2020; Gezari, 2021).
@@
Most TDEs that have been discovered so far exhibit a smooth single-peak light curve in both UV and optical bands. However, some of them display a bump in the rising phase of their light curves (Holoien et al., 2019; Charalampopoulos et al., 2023; Faris et al., 2023; Huang et al., 2023b; Wang et al., 2023b, Lin et al. in preparation.). This early signature may be produced by a number of physical processes, such as the cooling of unbounded debris, vertical compression during the first pericenter passage, shock breakout of self-crossing debris and so-called “TDE encore” (Kasen & Ramirez-Ruiz, 2010; Yalinewich et al., 2019; Wang et al., 2023b; Ryu et al., 2024). But no definitive conclusion has been reached. The collision between stream debris and the disk was proposed to explain the early bump in recent UV/optical outbursts of the candidate for repeated partial TDE (pTDE) ASASSN-14ko (Huang et al., 2023b). Furthermore, a fraction of the sources appeared to rebrighten in the optical light curve as the duration of observation increased (Jiang et al., 2019; Malyali et al., 2021; Chen et al., 2022; Wang et al., 2023c; Yao et al., 2023). Rebrightening in TDEs has also been found in X-rays, and the likely source is a repeated pTDE (Liu et al., 2023, 2024; Miniutti et al., 2023; Wevers et al., 2023). A repeated pTDE has been applied to explain the periodic optical outburst of ASASSN-14ko (Payne et al., 2021) and the optical rebrightening in AT 2020vdq (Somalwar et al., 2023).
@@
The source of the X-ray emission is unquestionably the accretion processes, but there is an ongoing debate regarding the origin of the UV/optical emissions. Two rival models are being discussed: one suggests that UV/optical emission is a result of the reprocessing of X-ray photons in an extended envelope/outflow (Loeb & Ulmer, 1997; Strubbe & Quataert, 2009; Metzger & Stone, 2016; Roth et al., 2016; Dai et al., 2018), while the other proposes that the emissions are caused by shocks generated by the stream-stream collision of the stellar debris (Piran et al., 2015; Jiang et al., 2016; Huang et al., 2023c; Jankovič et al., 2023). An intriguing phenomenon in optical TDEs is that X-ray emission lags UV/optical emission (Gezari et al., 2017; Pasham et al., 2017; Shu et al., 2020; van Velzen et al., 2021; Chen et al., 2022; Liu et al., 2022; Wang et al., 2022, 2023c; Huang et al., 2023a). This can be explained by a structural change in the reprocessing layer as the accretion rate declines from super-Eddington to sub-Eddington (e.g. Thomsen et al. 2022) or by a delayed onset of accretion after circularization.
@@
AT 2023lli (RA=22:57:39.470, DEC=+40:32:40.02) is a TDE discovered by the Gravitational-wave Optical Transient Observer (GOTO) on 2023 Jun 23 at a redshift of 0.036111https://www.wis-tns.org/object/2023lli. We conducted multiwavelength follow-up observations of this source and found that it showed a strong UV/optical bump and a delayed and episodic X-ray emission. In Section 2, we describe the observations and data analysis methods, and the multiwavelength results are represented in Section 3. In Section 4, we discuss the possible physical mechanisms for the early bump and the X-ray emission. In Section 5, we summarize our main results and conclusions. In this work, we assume cosmological parameters of , , and .
@@ 2.1 Swift X-Ray Photometry
We proposed several target-of-opportunity (ToO) observations to monitor the multiwavelength evolution using the X-Ray Telescope (XRT; Burrows et al. 2005) and the Ultraviolet/Optical Telescope (UVOT; Roming et al. 2005) on board the Swift Observatory. We also retrieved the public data (PI: Hinkle) from the High Energy Astrophysics Science Archive Research Center website. Swift data were processed with HEASoft 6.32.1. We ran the tasks xrtpipeline and xrtproducts to generate light curves and spectra. The source and background count rates were extracted from circular regions centered on the object with a radius of and a nearby circular region with a radius of , respectively. Source detection was performed using the ximage task. For data in which the source was not detected, we derived a 3 upper limit on the X-ray flux using the WebPIMMS tool222https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl and assuming a power-law X-ray spectrum with a photon index of 1.75 (Ricci et al., 2017). To improve the signal-to-noise ratio (S/N), we stacked the event files for MJD 60198-60201 (ObsIDs: 00016100029 and 00016100030) and MJD 60226-60244 (ObsIDs: 00016100038, 00016100039, and 00016100040), respectively. In these two stacked files, the source was detected above the 3 level using the ximage task.
@@ 2.2 XMM-Newton Observation
We proposed a ToO observation with XMM-Newton, which was performed on MJD 60295.5 (ObsID: 0932391301) with a total observation time of 20 ks. The XMM-Newton data were reduced with the Science Analysis System (version 21.0.0). The tasks cifbuild and odfingest were executed for data preparation, and then light curves and spectra were extracted through the task xmmextractor.
@@ 2.4 Ground-based Optical Photometry
AT 2023lli was first detected by the Asteroid Terrestrial impact Last Alert System (ATLAS; Tonry et al. 2018; Smith et al. 2020). The ATLAS - and -band light curves were obtained using the ATLAS Forced Photometry Service, which produces point spread function (PSF) photometry on the difference images. ATLAS has three to four single exposures within each epoch (typically within one day), so we binned the light curve every epoch to improve the S/N.
Upon noticing that AT 2023lli displayed a rebrightening after the initial peak (the “early bump”), we promptly conducted optical monitoring in the bands using 1.0 m telescopes from the Las Cumbres Observatory Global Telescope (LCOGT) network (Brown et al., 2013) starting from 2023 July 28. The observation continued until October 18, when it almost faded out in the LCOGT exposures. The observations were carried out every 2–3 days with a denser sampling around the peak, using exposure times of 260s, 160s, and 260s in the , , and bands, respectively.
Subsequently, we also initiated a follow-up campaign with the 2.5 m Wide Field Survey Telescope (WFST) in the bands with an exposure time of 30 s from 2023 September 14. There is a gap between November 21 and December 29 due to engineering maintenance and testing. The exposure time was changed to 60 s after the observations were resumed. Briefly, the WFST is a newly established photometric facility equipped with a 2.5 m diameter primary mirror, an active optics system, and a mosaic CCD camera with 0.73 gigapixels on the primary focal plane, allowing high-quality image capture over a 6.5 field of view (Wang et al., 2023a). It was installed near the summit of Saishiteng Mountain in northwestern China in the summer of 2023 and entered its commissioning stage in September following the official release of the first-light image of the Andromeda galaxy333https://wfst.ustc.edu.cn/news-and-meetings/news/20230917/20230917/. Follow-up observations with the WFST are particularly important when AT 2023lli becomes fainter, making its detection challenging with the 1m LCOGT network.
@@ 2.5 Spectral Observations and Data Reduction
Shortly after its discovery, the SCAT group (Tucker et al., 2022) acquired a spectrum of AT 2023lli utilizing the Supernova Integral Field Spectrograph, which was installed on the UH88 telescope (Hinkle, 2023). This spectrum was retrieved from the Transient Name Server (TNS) website555https://www.wis-tns.org/object/2023lli. We then acquired four more spectra using the Double Spectrograph (DBSP) instrument mounted on the Hale 200 inch telescope at Palomar Observatory (P200, Oke & Gunn 1982). For these observations, the D55 dichroic was adopted to split the incoming light longer than or shorter than 5500 Å into separate red and blue channels. A grism of 600 lines blazed at 3780 Å was used for the blue arm, and a grism of 316 lines blazed at 7150 Å was used for the red arm. The data were then reduced using the Python package Pypeit (Prochaska et al. 2020a, b), which can fully automatically implement the standard reduction procedure for long-slit spectroscopic observations. Furthermore, we obtained another spectrum using the Yunnan Faint Object Spectrograph and Camera (YFOSC) on board the LiJiang 2.4 m telescope (Fan et al., 2015; Wang et al., 2019) on 2024 January 2. All spectra are shown in Figure 3. Broad (10,000 km s-1) and probable Helium lines are detected when the source is bright. As the TDE emission significantly diminished, the spectrum (+147 and +1365 days) is dominated by host galaxy emission, showing deep Balmer absorption (see Figure 3), consistent with the types of galaxies that TDEs are known to preferentially inhabit (French et al., 2016).
@@ 3.1 Multiwavelength Light Curves
The multiwavelength light curves of AT 2023lli show a complex structure with an early bump and a main peak in UV/optical bands. The source was first detected by GOTO on MJD 60118, with a magnitude of 18.84 mag in the L band (Hinkle, 2023). We followed up the initial observation by ATLAS and found that the source had already been visible on MJD 60112.57 with a magnitude of mag in the c band. The source was then brightened to mag in the same band by MJD 60116.56, before fading. We missed the early UV emission of the source. The first UVOT observation was made on MJD 60125.5, when the source was already declining, with host-subtracted magnitudes of , and mag for the UVW1, UVM2 and UVW2 bands, respectively. The source then rose again after MJD 60140 and reached the optical peak on MJD 60164, with the magnitudes of , and mag for the g, r and i bands, respectively. At that time, we obtained magnitudes of , , , and mag for the UVW1, UVM2, UVW2 and U bands, respectively.
@@
We stacked a total of 27 Swift/XRT event files with an effective exposure time of 59.1 ks, covering the period from MJD 60125 to 60188. From these stacked data, we obtained a 3 upper limit of 0.0011 corresponding to a luminosity of by fixing the Galactic hydrogen column density value of (HI4PI Collaboration et al., 2016) and assuming a photon index of 1.75 (Ricci et al., 2017). The source was initially detected in the X-ray band until MJD 60198, where a 2.5 detection was achieved with a count rate of . Additionally, the source was also detected on MJD 60200. To improve the S/N, we combined the event files from two ObsIDs (00016100029 and 00016100030), which covered the time period from MJD 60198 to 60200. From these stacked data, we derived a mean count rate of at the 3 level, with an effective exposure time of 4013 s. However, in a subsequent observation with Swift/XRT lasting 4152 s, only a 3 upper limit of 0.004 was obtained. By merging three event files from MJD 60226 to 60237 (ObsIDs: 00016100038–00016100040), we derived a mean count rate of . In the following Swift/XRT observation with an exposure time of 3678 s, the source was not detected, and a 3 upper limit of 0.003 is imposed. Thus, the X-ray emission from the source appeared to exhibit a sporadic behavior with a general upward trend. At MJD 60295.5, XMM-Newton recorded an unabsorbed luminosity of in the X-ray. Following this, the X-ray emission was detectable in every subsequent observation with XRT. When the X-ray showed sporadic behavior, we checked the data and found that the count rate was above the statistical fluctuation and confirmed that such behavior actually existed (P-value of 0.0004). For XRT, the conversion factor from count rate to flux was determined to be , which was derived from the spectral fit of the stacked spectrum extracted from all event files with X-ray detection.
@@
The UV/optical SEDs were fitted with a blackbody model and SuperBol (Nicholl, 2018) was used to calculate the temporal evolution of the blackbody luminosity (), temperature () and blackbody radius (). The results are displayed in Figure 4. The highest was recorded on MJD 60164.6 with a value of and at that time we derived and . After that, the luminosity decreases, while no significant temporal evolution can be seen in the blackbody temperature. We derived the energy released in the UV/optical bands of and during the bump and the main outburst, respectively. Using the emcee python package (Foreman-Mackey et al., 2013), we employed a “Gaussian rise and power-law decay” model to fit in both the bump and the main outburst, as described in van Velzen et al. (2019, 2021); Hammerstein et al. (2023); Yao et al. (2023). We obtained a power-law index of -4.10, which is much steeper than the theoretical prediction of -5/3 (Rees, 1988). Moreover, some researchers found that TDEs exhibit a wide range of power-law indices, from -4.0 to -0.5 (van Velzen et al., 2019, 2021; Gezari, 2021; Hammerstein et al., 2023). Our result resembles that of AT 2019mha, the most rapidly fading source (with a power-law index of 4.0) among the optically selected TDEs studied by van Velzen et al. (2021). This makes AT 2023lli one of the fastest declining TDEs ever observed. A steep slope in the light curve has been predicted by some theoretical work (Guillochon & Ramirez-Ruiz, 2013; Ryu et al., 2020) that involves pTDEs. As showen by Ryu et al. (2020), a more rapid decrease in light curves is associated with a lower amount of mass being removed from the star. Consequently, the situation of AT 2023lli could be connected to a pTDE. A recent systematic analysis conducted on optically-selected TDEs has found that the ratio between UV/optical and X-ray luminosities () tends to approach values similar to those of a disk () at later times, although it exhibits a wide range of values at earlier times (Guolo et al., 2023). This trend is also observed in the case of AT 2023lli. As the X-ray emission becomes visible and increases in brightness, the ratio gradually increases, with the most recent value slightly above 1 (refer to the bottom panel of Figure 4).